# Universal coarse geometry of spin systems

### v1

@@ 7.1. Topological order: superselection sectors
To prove the second point, suppose $(X,d)=\mathbbm{Z}^{n}$ with the restriction of the Euclidean metric. For any $K,B$, we note that is some $R_{K,B}\geq 0$ such that for any coarse $K,B$ cone $\mathcal{D}$, there is a $1,0$ cone $\mathcal{D}^{\prime}$ with $\mathcal{D}^{\prime}\subseteq\mathcal{D}{+R_{K,B}}$. Thus if $H\in\text{Sec}_{\phi}$, then $H$ is localizable in $\mathcal{D}^{\prime}$, then it is localizble in $\mathcal{D}+R_{K,B}$ so $H\in\text{Sec}^{d}_{\phi}.$
Now let $H\in\text{Sec}^{d}_{\phi}$, and $\mathcal{D}$ and honest cone. Choose a cone $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ such that $\mathcal{D}^{\prime}+R_{1,0}\subseteq\mathcal{D}$. Then $H$ is localized in $\mathcal{D}$, hence $H\in\text{Sec}_{\phi}$.
In the case when $(X,d)$ is coarsely equivalent to $\mathbbm{Z}^{2}$, this category can be upgraded with the structure of a braided W*-tensor under a condition called approximate Haag duality [Oga22], generalizing the familiar Haag duality from AQFT utilized in earlier approaches [Naa11, Naa12, FN15, CNN18]. It is a bit tedious but not difficult to check using the style of argument from our proposition that this condition can be reformulated in terms of coarse cones, and that the construction of the tensor product and braiding can be carried out in the coarse setting. Alternatively, one could show that an apropriate poset of coarse cones satisfies the axioms in [BBC+24], which by their main reults would allow a direct definition of braided tensor category structure on $\text{Sec}^{d}_{\phi}$.

### v2

@@ 7.1. Topological order: superselection sectors
To prove the second point, suppose $(X,d)=\mathbbm{Z}^{n}$ with the restriction of the Euclidean metric. Note that for any $K,B,\mathcal{D}$, there is some $R\geq 0$ such that the $R$-neighborhood of the image of any $K,B,\mathcal{D}$ coarse cone contains an honest Euclidean cone (i.e. the intersection of a cone in $\mathbbm{R}^{n}$ with the lattice $\mathbbm{Z}^{n}$). Now let $f:\mathcal{D}\rightarrow\mathbbm{Z}^{n}$ be a coarse $K,B,\mathcal{D}$ cone. If $H\in\text{Sec}_{\phi}$, then since $H$ is localizable in $\mathcal{D}^{\prime}$, since $\mathcal{D}^{\prime}\subseteq f(\mathcal{D})+R$, $H$ is also localizble in $f(\mathcal{D})+R$. Since $K,B,\mathcal{D}$ and $f$ were arbitrary, $H\in\text{Sec}^{d}_{\phi}.$
Now let $H\in\text{Sec}^{d}_{\phi}$, and $\mathcal{D}$ an honest cone. If $\mathcal{D}=\mathbbm{Z}^{n}\cap\mathcal{D}^{\prime}$ for $\mathcal{D}^{\prime}$ a cone in $\mathbbm{R}^{n}$, then $\mathcal{D}$ is $1,\sqrt{2},\mathcal{D}^{\prime}$ cones, where $f:\mathcal{D}^{\prime}\rightarrow\mathcal{D}$ is the map that chooses the nearest element in $\mathcal{D}$. Note that for any vector $v\in\mathbbm{Z}^{n}$, the honest cone $\mathcal{D}^{\prime\prime}=\mathcal{D}+v$ is also a $1,\sqrt{2},\mathcal{D}^{\prime}$ coarse cone. Let $R$ be the $1,\sqrt{2},\mathcal{D}^{\prime}$ localizing constant. Then by translating $\mathcal{D}$ into its interior, we can find an honest cone $\mathcal{D}^{\prime\prime}$ such that $\mathcal{D}^{\prime\prime}+R\subseteq\mathcal{D}$. Then $H$ is localized in $\mathcal{D}^{\prime\prime}\subseteq\mathcal{D}$ hence also in $\mathcal{D}$, and thus $H\in\text{Sec}_{\phi}$.
In the case when $(X,d)$ is coarsely equivalent to $\mathbbm{Z}^{2}$, this category can be upgraded with the structure of a braided W*-tensor category under a condition called approximate Haag duality [Oga22], generalizing the familiar Haag duality from AQFT utilized in earlier approaches [Naa11, Naa12, FN15, CNN18]. It is a bit tedious but not difficult to check using the style of argument from our proposition that this condition can be reformulated in terms of coarse cones, and that the construction of the tensor product and braiding can be carried out in the coarse setting. Alternatively, one could show that an apropriate poset of coarse cones satisfies the axioms in [BBC+24], which by their main reults would allow a direct definition of braided tensor category structure on $\text{Sec}^{d}_{\phi}$.